OpenStudy (anonymous):

i'm not sure how to do this Calculate 1 Liter of 0.1M hydrochloric acid (MW=36.6). HCl has a purity of 37% and a density of 1.19g/ml

OpenStudy (photon336):

I'll gladly help. What are you being asked to do? it's a bit unclear.

OpenStudy (anonymous):

Hi, thank you! this is exactly how it is, I couldn't understand it nor how to start with it.

OpenStudy (anonymous):

Do you have any idea how to do this?

OpenStudy (photon336):

Well here's what I think. Take a look at each of these steps and ask if you have questions. For starters, molarity means moles per liter 1. so 0.1 M means that there are 0.1 moles of HCl per liter. 2. \[0.1M = \frac{ 0.1 moles~HCl }{ L }\] 3. Now in one liter of HCL. to find the number of moles contained in our solution we must multiply the molarity by liters. I'll show you this see how liters cancel out and we're left with 0.1 moles. 4. \[1~liter*(0.1)\frac{ mol }{ liter } = 0.1~mol~HCl\] 5. so our sample has 0.1 moles of HCL we're given the molar weight of HCL which is 37 grams. so every mole of HCL contains 37 grams. we know that we have 0.1 mol of HCL, 7. \[0.1~moles*(\frac{ 37grams }{ mole }) = 3.7~grams~of~HCl\] so 0.1 M sample of HCL contains 3.7 grams of HCl. in 1 liter of solution.

OpenStudy (anonymous):

makes sense, thank you for explaining :)

OpenStudy (anonymous):

So 36% and 1.19g/ml are just there to confuse me? lol

OpenStudy (photon336):

yeah those are strange let me explain this.

OpenStudy (photon336):

thing is you already know what the molarity is and how much volume in Liters is in your sample right? so you kind of don't need density. If I understood this correctly. there's a lot of information that you can figure out with the information. Also, purity. means that in HCl or I guess this sample of HCl there are impurities. i'm wondering though was this from a lab experiment?

OpenStudy (anonymous):

No, this was a practice problem for tomorrow's cell bio lab quiz. We didn't have time to go over it in class.

OpenStudy (photon336):

Thing is you know what the molarity is right? and you know the volume it's given to you as 1 liter. so you can easily find the number of moles and grams of HCL are in that sample. the purity just tells you, hey my sample is impure, I need to fix that, in lab.

OpenStudy (photon336):

say if I had 0.5 Liters of solution, and the molarity was 0.1 M how many grams of HCL would be in my solution?

OpenStudy (anonymous):

18.23 g HCl 

OpenStudy (anonymous):

right?

OpenStudy (photon336):

almost \[0.5~liters*(\frac{ 0.1~moles }{ L }) = 0.5*0.1 =0.05~moles~HCL\]

OpenStudy (photon336):

take a look at what I did above

OpenStudy (anonymous):

.05 sorry x 36.46 = 1.82

OpenStudy (anonymous):

I get it now, thank you for your time! :)

OpenStudy (photon336):

Anytime, good luck.

OpenStudy (anonymous):

thanks